OpenStudy (anonymous):

Find the term that should be added to the expression to form a perfect square trinomial. Write the resulting perfect square trinomial in factored form. x^2+11x

OpenStudy (anonymous):

well, we have \((a+b)^2 = a^2 + 2ab + b^2\), so by analogy, what could \(a\) and \(b\) be?

OpenStudy (anonymous):

11/2)^@= 121/4= x^2+11x+121/4= (x+11/2)^2 but it's saying that's incorrect.

OpenStudy (anonymous):

That is undoubtedly a correct answer. I am unsure why your book says otherwise. Technically, you could also add on \(-11x\), making for \((x+0)^2\).

OpenStudy (anonymous):

What is the constant term of the equation is the 121/ 4?

OpenStudy (anonymous):

Got it, it was. whoop.

OpenStudy (anonymous):

x^2 + 11x + 121/4 (x + 11/2)(x + 11/2) that's ur answer :)

OpenStudy (anonymous):

never done a question even remotely close too this lol

OpenStudy (anonymous):

Pff, I never even done this when I was in high school!

OpenStudy (anonymous):

lol im in high school and i can confirm ive never done this lol but im aboout too leave :D cannt wait!